# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v2

@@ VII-B Benchmark Schemes
Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@5050{50}50 (s)
@@
  • Note: $30$+ or $1200$+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
VII-C Performance Evaluation of MDDR
First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach. We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig. 7 and Table II. We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR. This is because MAPPO is a centrally controlled distributed execution [44] that can utilize more information for decision making. However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs. For the fully distributed MDDR, it is more suitable for non-cooperative scenarios. The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared. This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly. Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time. The behaviors of each MU and MSP in the trading are shown in Figs. 8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
VII-D Benchmark Comparison
@@
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing learning models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training. As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace. In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased. Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
@@
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because the potential value $\omega$ is higher for MSP 3, leading to more rewards being given by MSP 3. As a result, the MUs will allocate more computational and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource. Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Yet, the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them. The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU. Additionally,from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $\omega$ associated with MSP 2 and MSP 3.
@@
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP@5050{50}50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Meanwhile, Figs. 10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively.
@@ VIII Conclusion
To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to contribute locally trained data for AR services such as object detection and classification. Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests. A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the learning models of MUs for AR services. In addition, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs. Extensive simulations on AR-related vehicle and MNIST datasets demonstrate that the proposed framework enables more efficient and immersive AR services in the vehicular metaverse.
@@ Appendix A Proof of Proposition 1
$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (24)
$-\sum_{n}\left[c_{m}^{f}f_{mn}\log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
@@
$-\sum_{n}\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
@@
where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively. Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
@@
$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}.$
From (25) and the constraints in (11), the Karush-Kuhn-Tucker (KKT) [47] conditions are given by
@@
$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
@@
$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
@@
Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
@@
  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
@@
    $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}.$
@@
Appendix B Proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$
@@
$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
@@
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
@@
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
@@
Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40). Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$. For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (22) and obtain the following result as shown in (41). Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved.
@@
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}\log(1/\theta_{m}),a_{2}=\log_{2}(1+\varsigma_{mn}).$

### v3

@@ VII-D Benchmark Comparison
Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@50{50} (s)
@@
  • Note: 3030+ or 12001200+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
@@
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing local models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the local model but also considers both the amount and potential value of raw data used for training. As a result, resource-constrained MUs are motivated to contribute more resources, providing valuable local models to MSPs more rapidly.
In contrast, baseline schemes like x_based, w_based, and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if an MU has already reached its maximum allocatable resources at a given reward $p$, increasing $p$ further will not yield additional contributions. In this situation, raising the reward only increases the MSP’s cost without eliciting additional contributions, so its utility actually declines. Consequently, the uniform reward $p$ remains unchanged, preventing other MUs with more resources from contributing higher-value local models.
@@
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because MSP 3 is associated with a higher potential value $\omega$, resulting in more substantial rewards being offered. As a result, the MUs will allocate more computational and communication resources to provide local models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). Notably, the fixed scheme performs the worst due to the limitations of synchronous FL, which is bottlenecked by the least capable MU.
Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Although the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based, its performance improvement rate is comparatively slower. This is because the IoM represents the aggregate value across all MUs, while the performance improvement rate is hampered by the worst MU. Furthermore, from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $\omega$ associated with MSP 2 and MSP 3.
@@
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table IV gives the mAP@50{50} (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. Compared to MNIST, the performance enhancement for object detection is slower due to the complexity of the detection task. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Moreover, Figs. 10(d)10(e)10(f) and Table IV demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively.
@@ VIII Conclusion
To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework, incorporating FL to incentivize MUs to contribute local models for AR services such as object detection and classification. Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests. A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the local models of MUs for AR services. Furthermore, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs. Extensive simulations on AR-related vehicle and MNIST datasets demonstrate that the proposed framework enables more efficient and immersive AR services in the vehicular metaverse.
@@ Appendix A Proof of Proposition 1
$\ L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)$ (24)
$=$ $\ \sum_{n}p_{mn}V_{mn}-\sum_{n}\left[c_{m}^{f}f_{mn}\log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
@@
$-\sum_{n}\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]$
@@
where ($\lambda_{m}\geq 0$ and $\beta_{m}\geq 0$), ($\delta_{mn}\geq 0$), and ($\gamma_{m}\geq 0$) are the Lagrangian dual variables corresponding to constraints C1, C2, and C3, respectively. Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
@@
$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})},$
respectively. From (25) and the constraints in (11), the Karush-Kuhn-Tucker (KKT) [47] conditions are given by
@@
$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]=0,$
@@
$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]=0,$
@@
Then, based on the value of $\delta_{mn}$, the optimal allocation of computational and communication resources exists in the following two cases:
@@
  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}\right]=0$, we can get
@@
    $F=2\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{2\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}.$
@@
Appendix B Proof of Theorem 2
@@
$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0,$ (34)
@@
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}}.$ (35)
@@
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}}.$ (38)
@@
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\Bigg[\frac{aI_{mn}^{3}p_{mn}+4aI_{mn}^{2}\delta_{mn}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{bI_{mn}^{3}p_{mn}+4bI_{mn}^{2}\delta_{mn}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{\mu_{n}I_{mn}^{2}\left(\frac{aI_{mn}}{2f_{mn}(p_{mn}I_{mn}+\delta_{mn})}+\frac{bI_{mn}}{2B_{mn}(p_{mn}I_{mn}+\delta_{mn})}\right)^{2}}{\left(1+\sum_{m}V_{mn}\right)^{2}}$ (41)
$\qquad\qquad\qquad+\frac{\mu_{n}I_{mn}\left(\frac{3aI_{mn}^{2}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{3bI_{mn}^{2}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}\right)}{\left(1+\sum_{m}V_{mn}\right)}\Bigg]<0,$
$\text{where}\,a=x_{m}\tau_{n}\log(1/\theta_{m}),\ b=\frac{b_{mn}}{\log_{2}(1+\varsigma_{mn})}.$
Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40). Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$. For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, we substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (22) and obtain the following result as shown in (LABEL:result). Next, by randomly selecting a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 0, we obtain
$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
$=$ $\underbrace{\sum_{m=1}^{M}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}}_{\text{diagonal part}}+\underbrace{\sum_{m\neq m^{\prime}}h^{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{{m^{\prime}}n}}h^{m^{\prime}}}_{\text{off-diagonal part}}.$ (42)
Expanding the expressions and letting $S=\sum_{m=1}V_{mn}$, we have
$\ \quad\bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
$=\sum_{m\neq{m^{\prime}}}h^{m}\left(-\mu_{n}\frac{V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}}{(1+S)^{2}}\right)h^{m^{\prime}}$
$+\sum_{m=1}^{M}\left[\mu_{n}\frac{V_{mn}^{\prime\prime}(1+S)-(V_{mn}^{\prime})^{2}}{(1+S)^{2}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right](h^{m})^{2}.$ (43)
Specially, the off-diagonal part is transformed using the identity
$\ \sum_{m\neq m^{\prime}}h^{m}V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}h^{m^{\prime}}$
$=$ $\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2},$ (44)
thus,
$\sum_{m\neq{m^{\prime}}}h^{m}\left(-\mu_{n}\frac{V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}}{(1+S)^{2}}\right)h^{m^{\prime}}$
$=$ $-\frac{\mu_{n}}{(1+S)^{2}}\left[\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2}\right].$ (45)
Then, combining both parts, we have
$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
$=$ $\sum_{m=1}^{M}\left[\mu_{n}\frac{V_{mn}^{\prime\prime}(1+S)-(V_{mn}^{\prime})^{2}}{(1+S)^{2}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right](h^{m})^{2}$
$\quad-\frac{\mu_{n}}{(1+S)^{2}}\left[\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2}\right].$ (46)
Moreover, the expression can be simplified to
$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
$=$ $-\frac{\mu_{n}}{(1+S)^{2}}\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}$
$+\sum_{m=1}^{M}(h^{m})^{2}\left(\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right).$ (47)
Finally, we analyze the sign of this expression:
  1. 1.
    Since $\mu_{n}>0$ and $(1+S)^{2}>0$, the first term is non-positive.
  2. 2.
    Plugging (40) into the second term yields $(\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime})<0$, as shown in (LABEL:second-term).
    $\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}=-\Bigg[\frac{aI_{mn}^{3}p_{mn}+4aI_{mn}^{2}\delta_{mn}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{\mu_{n}I_{mn}\left(\frac{3aI_{mn}^{2}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{3bI_{mn}^{2}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}\right)}{\left(1+S\right)}\Bigg]<0,$ (48)
    $\text{where}\,a=x_{m}\tau_{n}\log(1/\theta_{m}),\ b=\frac{b_{mn}}{\log_{2}(1+\varsigma_{mn})}.$
Based on the above, we can have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. According to [39], there exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$. ∎